# Graphical Calculus for Non-Gaussian Quantum States

### v1

@@ Abstract
Introduction.— In the quest to demonstrate a genuine quantum advantage in the continuous variable (CV) domain, substantial progress on the controllable generation of multimode non-Gaussian states has recently been made by photon subtraction  [1, 2, 3], nonlinear unitary operations, such as spontaneous parametric down- conversion [4] or Kerr interactions in microwave cavities coupled superconducting artificial atoms [5, 6]. Such strategies usually introduce higher-order statistical moments of the quadrature field operators, go beyond Gaussian states and therefore have irreplaceable advantages as quantum resource [7, 8] in various tasks, such as entanglement distillation [9, 10, 11], quantum error correction [12, 13], quantum metrology [6, 14], and CV quantum computation [15, 16, 17]. Compared to the remarkable progress in experimental realizations, how to understand and more definitely classify the multimode non-Gaussian states with high-order interplay among quadratures, still faces a formidable challenge. To address it in a more intuitive and systematic manner, the adoption of graphical formalisms [18, 19, 20, 21, 22, 23, 24] as a means of describing quantum states emerges as a highly promising solution.
@@
Graph states [25, 26, 27] and hypergraph states [28, 29, 30, 31, 32] were first defined in discrete-variable (DV) systems to form multiparticle entangled states corresponding to mathematical graphs. They have applications in various contexts, ranging from quantum error correction [33, 34] to measurement-based quantum computation [35, 36]. This kind of definition in terms of graphical formalism then got generalised to CV system [37, 38], and examples are shown in Fig. 1(a) and Fig. 1(b). It was proved that any Gaussian state can be described by a CV graph state [39], as the first and second-order statistical moments of the quadratures which completely characterise a Gaussian state can be derived from the edges in the corresponding graph. Naturally, CV hypergraph states have richer graphical structures [40, 41], since they include multimode interactions. Such inherent multimode nonlinearity can be used to represent the high-order correlation in non-Gaussian states. Ref. [40] showed an evidence that a hypergraph state restricted to a Gaussian measurement strategy can be a candidate for universal quantum computing. Nevertheless, the ramification of arbitrary Gaussian operations on non-Gaussian hypergraph states remains an underexplored frontier, necessitating further investigation.
@@
In this paper, we provide a unified graphical calculus that serves as a comprehensive framework for manipulating non-Gaussian states in CV systems. This calculus encapsulates hypergraph transformation rules encompassing all single-mode and typical two-mode Gaussian unitary operations, together with local quadrature measurements, offering an intuitive visual modality for implementing experimentally feasible operations on CV non-Gaussian states. Crucially, as shown in Fig. 1(c), we demonstrate the capability to generate complex hypergraph states with more or higher-order hyperedges from simple structures solely through Gaussian operations, underscoring the power and versatility of the graphical rules provided here. Furthermore, we give three examples of non-Gaussian state preparation schemes rooted in our graph-based formalism, which hold promise for application in the broader landscape of CV universal quantum computing. These instances illustrate the practical potential of the graphical calculus, positioning it as a valuable tool for advancing the frontiers of quantum information processing.
@@
CV Hypergraph States.— A weighted hypergraph is a tuple $H=(V,E,T)$ with a set $V=\{v_{i}\}_{i=1}^{n}$ of $n$ vertices, a set $E=\{e_{j}\}_{j=1}^{|E|}$ of ${|E|}$ hyperedges, and a set $T=\{t_{e_{j}}\}_{j=1}^{|E|}$ of real valued weights associated to hyperedges, where $|E|$ denotes the number of elements of $E$. A CV hypergraph state |$\ket{H}$ is defined as an $n$-mode quantum state where the vertices represent the modes of the state and the weighted hyperedges represent generalized controlled-$Z$ gates performed on the modes [41, 40]. More formally, we have |$\ket{H}=\prod_{e_{j}\in E}C_{e_{j}}(t_{e_{j}})\ket{0}_{p}^{\otimes n}$, where |$\ket{0}_{p}$ is the zero-momentum eigenstate in the limit of infinite squeezing with normalization $\bra{p^{\prime}}\ket{p}=\delta_{p,p^{\prime}}$ and $C_{e_{j}}(t_{e_{j}})=\operatorname{e}^{it_{e_{j}}\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}}$ is a generalized controlled-$Z$ gate acting on $m$ modes connected by the $j-$th hyperedge $e_{j}=\{v_{1},v_{2},\dots v_{m}\}$. If a hyperedge only contains two vertices, i.e. $\lvert e\rvert=2$, $C_{e}$ is the standard controlled-$Z$ gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}}$ and if it contains three vertices, it is the Toffoli gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\operatorname{\mathit{\hat{q}}}_{v_{3}}}$ [42]. We use the shorthand notation $q_{e_{j}}\coloneqq\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}$. A hypergraph state of $k$-th order is classified by the maximum number of $k$ modes connected with one hyperedge.
@@
For clarity, we give an example of a 4-mode hypergraph state containing $|E|=2$ hyperedges, as shown in Fig. 2(a). One hyperedge $e_{1}=\{A,B,C\}$ has weight $t_{e_{1}}=1$ and the other hyperedge $e_{2}=\{A,D\}$ has weight $t_{e_{2}}=2$. This hypergraph state is given by
|$\ket{H}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{D}}\ket{0}_{p}^{\otimes 4}.$ (1)
@@
Figure 2: Examples of a weighted hypergraph state and different Gaussian operations applied on it. Roman labels represent the weight of each hyperedge. The orange circle represents the initial hyperedge, and the green circle represents the newly created hyperedge. The purple solid lines and circles represent the 2nd-order and 1st-order hyperedges respectively. The dotted circles represent the higher-order moments. (a) A hypergraph $H$, corresponding to four modes and two hyperedges: one 2nd-order hyperedge $\{A,D\}$ with weight 2 and the other 3rd-order hyperedge $\{A,B,C\}$ with weight 1. On the corresponding quantum state, one can apply different Gaussian operations: (b) Applying momentum displacement $Z^{B}(-2)$ on mode $B$ adds an hyperedge representing $\operatorname{e}^{-i2\operatorname{\mathit{\hat{q}}}_{B}}$ on it. (c) Applying position displacement $X^{A}(1)$ on mode $A$ generates new hyperedges: a 1st-order hyperedge representing the term $\operatorname{e}^{-i2\operatorname{\mathit{\hat{q}}}_{D}}$ on $D$ and a 2nd-order hyperedge representing $\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$ for $\{B,C\}$. (d) Applying position shearing on mode $B$, $D_{q}^{B}(2)$, adds a decoration representing $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}}$ on $B$, which is not part of the formalism. (e) Similarly, momentum shearing $D_{p}^{A}(1)$ on mode $A$ adds decorations, representing $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$ and $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, which fall outside the formalism on $D$ and $B,C$, but also induces a valid new hyperedge representing $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$ for $\{B,C,D\}$. (f) Squeezing $S^{A}(1)$ on mode $A$ changes the weights on all hyperedges adjacent to mode $A$. (g) Rotation $R^{D}(\pi)$ on mode $D$ changes the weights by inverting all signs on hyperedges adjacent to $D$. (h) Controlled-$Z$ gate $C_{A,D}(0.5)$ on mode $A$ and mode $D$ adjusts their weight by 0.5.
@@
In order to present the graphical calculus more conveniently, we first introduce some basic rules of set operations. The hyperedge set $E_{a}$ is defined as the set of all hyperedges containing mode $a$: $E_{a}=\{e\mid e\in E,a\in e\}$. Furthermore, the adjacency set $\operatorname{\mathcal{A}}(a)$ of vertex $a\in V$ is defined as $\operatorname{\mathcal{A}}(a)=\{e\setminus\{a\}\mid e\in E_{a}\}$. To consider the bijective mapping between hyperedge and weight sets, we define the union $(E_{U};T_{U})=(E_{1};T_{1})\cup(E_{2};T_{2})$ as a union of all tuples. For hyperedges appearing in both sets, we add up their weights. More precisely, for any $(e,t_{1})\in(E_{1};T_{1})$ and $(e,t_{2})\in(E_{2};T_{2})$, we get $(e,t_{1}+t_{2})\in(E_{U};T_{U})$. For the hypergraph in Fig. 2(a), the set of hyperedges containing mode $A$ is given by $E_{A}=\{\{A,B,C\},\{A,D\}\}$ with associated weights $T_{A}=\{1,2\}$ and $\mathcal{A}(A)=\{\{B,C\},\{D\}\}$. If we want to unite $E_{A}$ and $E_{B}=\{\{A,B,C\}\}$ with $T_{B}=\{1\}$, we obtain $(E_{U};T_{U})=(\{\{A,B,C\},\{A,D\}\};\{2,2\}).$
@@
Although not all quantum states can be transformed under the graphical rules, some states are able to be represented indirectly in the form of hypergraphs, for example the cubic phase state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}\ket{0}_{p}$ which contains a third-order moment term of one single-mode operator and plays important roles in universal quantum computing [43, 15]. It turns out that it can be prepared in our formalism.
@@
Gaussian Operations on CV Hypergraph States.— In this part we provide unified graphical rules of Gaussian unitary operations and Gaussian measurements for CV hypergraph states. We show the graphical rules of single-mode Gaussian operations including displacement, shearing, squeezing, and rotation [44, 45] and further discuss the CV controlled-$Z$ gate. Derivations are given in the appendix. As any multimode Gaussian operation can be represented as a concatenation of a beam splitter, single-mode squeezing, and phase shifts through the Bloch-Messiah decomposition [46], and the beam-splitter operation can be replaced by controlled-$Z$ gate operations [42], any multimode Gaussian operation can be represented in our graphical calculus.
@@
Operation I: Displacement.— The displacement operators include the momentum displacement $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ and the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$, where $s$ represents the distance that the operator $\operatorname{\mathit{\hat{q}}}_{a}$ or $\operatorname{\mathit{\hat{p}}}_{a}$ is moved in phase space [47]. Thus, the momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ on mode $a$ modifies the hyperedge set by $(E^{\prime};T^{\prime})=(E;T)\cup(\{a\};s)$. In other words, if mode $a$ has its own hyperedge initially, this operation will increase the weight by $s$. If not, a new hyperedge with the weight $s$ will be added to the system. Fig. 2(b) shows an example of performing operation $Z^{B}(-2)$ on the hypergraph state depicted in Fig. 2(a).
@@
However, when performing the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$ to the hypergraph state, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, additional correlation terms will be generated when they simultaneously act on the zero-momentum eigenstate |$\ket{0}_{p}$. It influences hyperedges of the adjacency $\mathcal{A}(a)$ of mode $a$ such that $(E^{\prime},T^{\prime})=(E,T)\cup(\operatorname{\mathcal{A}}(a),\{\tilde{t}\})$, where $\tilde{t}\coloneqq-s\times t_{e\cup a}$, which means the position displacement on mode $a$ will add or change hyperedges between the adjacent modes to mode $a$ with the weight multiplied by $-s$. For example, when the position displacement operator $X^{A}(1)$ is applied to mode $A$ of the hypergraph state corresponding to Fig. 2(a), two new hyperedges between vertices $\{B,C\}$ and $\{D\}$ of $\operatorname{\mathcal{A}}(a)$ are created, as represented by the purple line and purple circle in Fig. 2(c).
@@
From the above analysis, it can be seen that in a hypergraph state of $k$-th order, displacement operations can create new hyperedges of at most order $k-1$, which is similar to rules holding for multi-qubit hypergraph states [31, 48]. Note that for CV graph states, the order of edges is $k=2$ and therefore displacement operations only cause local effects in a certain mode.

### v2

@@ Abstract
1 Introduction
In the quest to demonstrate a genuine quantum advantage in the continuous variable (CV) domain, substantial progress on the controllable generation of multimode non-Gaussian states has recently been made by photon subtraction [1, 2, 3] and nonlinear unitary operations such as spontaneous parametric down-conversion (SPDC) [4]. These strategies usually introduce higher-order statistical moments of the quadrature field operators, go beyond Gaussian states and therefore have irreplaceable advantages as quantum resource [5, 6] in various tasks, including entanglement distillation [7, 8], quantum metrology [9], and CV quantum computation [10, 11, 12, 13]. Compared to the remarkable progress in experimental realizations, how to understand and more definitely classify the multimode non-Gaussian states with high-order interplay among quadratures, still faces a formidable challenge. To address it in a more intuitive and systematic manner, graphical formalisms [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] as a means of describing quantum states emerges as a highly promising solution.
@@ 1 Introduction
Graph states [24, 25, 26] and hypergraph states [27, 28, 29, 30, 31] were first defined in discrete variable (DV) systems to form multiparticle entangled states corresponding to mathematical graphs. They have applications in various contexts, ranging from quantum error correction [32, 33] to measurement-based quantum computation [34, 35]. This kind of definition in terms of graphical formalism then got generalised to CV system [36, 37]. Examples are shown in Figs. 1(a) and (b). It has been proven that any Gaussian state can be described by a CV graph state [38]. The first and second-order statistical moments of the quadratures, which completely characterise a Gaussian state, can be derived from the edges in the corresponding graph. Naturally, CV hypergraph states have richer graphical structures [39, 40], since they include multimode interactions. Such inherent multimode nonlinearity can represent the high-order correlation in non-Gaussian states. Ref. [39] showed an evidence that a 3-order hypergraph state restricted to a Gaussian measurement strategy can be a candidate for universal quantum computing. Nevertheless, the ramification of arbitrary Gaussian operations on non-Gaussian hypergraph states remains an underexplored frontier, given their potential to yield higher-order hypergraph states that distribute non-Gaussian resources across an expanded set of modes.
@@
Here, we provide a unified graphical calculus that serves as a comprehensive framework for manipulating non-Gaussian states in CV systems. This calculus encapsulates hypergraph transformation rules encompassing all single-mode and typical two-mode Gaussian unitary operations, together with local quadrature measurements, offering an intuitive visual modality for implementing experimentally feasible operations on CV non-Gaussian states. Crucially, as shown in Fig. 1(c), we demonstrate the capability to generate complex hypergraph states with more or higher-order hyperedges from simple structures solely through Gaussian operations, underscoring the power and versatility of the graphical rules provided here. Furthermore, we give three examples of non-Gaussian state preparation schemes rooted in our graph-based formalism. These instances illustrate the practical potential of the graphical calculus, positioning it as a valuable tool for advancing the frontiers of quantum information processing.
2 CV Hypergraph States
@@
A weighted hypergraph is a tuple $H=(V,E,T)$ with a set $V=\{v_{i}\}_{i=1}^{n}$ of $n$ vertices, a set $E=\{e_{j}\}_{j=1}^{|E|}$ of ${|E|}$ hyperedges, and a set $T=\{t_{e_{j}}\}_{j=1}^{|E|}$ of real valued weights associated to hyperedges, where $|E|$ denotes the number of elements of $E$. A CV hypergraph state |$\ket{H}$ is defined as an $n$-mode quantum state where the vertices represent the modes of the state and the weighted hyperedges represent generalized controlled-$Z$ gates performed on the modes [40, 39]. More formally, we have |$\ket{H}=\prod_{e_{j}\in E}C_{e_{j}}(t_{e_{j}})\ket{0}_{p}^{\otimes n}$, where |$\ket{0}_{p}$ is the zero-momentum eigenstate in the limit of infinite squeezing with normalization $\bra{p^{\prime}}\ket{p}=\delta_{p,p^{\prime}}$ and $C_{e_{j}}(t_{e_{j}})=\operatorname{e}^{it_{e_{j}}\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}}$ is a generalized controlled-$Z$ gate acting on $m$ modes connected by the $j-$th hyperedge $e_{j}=\{v_{1},v_{2},\dots v_{m}\}$. If a hyperedge only contains two vertices, $C_{e}$ is the standard controlled-$Z$ gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}}$ and if it contains three vertices, it is the Toffoli gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\operatorname{\mathit{\hat{q}}}_{v_{3}}}$ [41]. We use the shorthand notation $q_{e_{j}}\coloneqq\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}$. A hypergraph state of $k$-th order is classified by the maximum number of $k$ modes connected with one hyperedge.
@@
For clarity, we give an example of a four-mode hypergraph state containing two hyperedges, |$\ket{H}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{D}}\ket{0}_{p}^{\otimes 4}$, as shown in Fig. 2(a). One hyperedge $e_{1}=\{A,B,C\}$ has weight $t_{e_{1}}=1$ and the other hyperedge $e_{2}=\{A,D\}$ has weight $t_{e_{2}}=2$.
@@
Figure 2: Examples of a weighted hypergraph state and different Gaussian operations applied on it. Roman labels represent the weight of each hyperedge. The orange circle represents the initial hyperedge, and the green circle represents the newly created hyperedge. The purple solid lines and circles represent the 2nd-order and 1st-order hyperedges respectively. The dotted circles represent the higher-order moments. (a) A hypergraph $H$, corresponding to four modes and two hyperedges: one 2nd-order hyperedge $\{A,D\}$ with weight 2 and the other 3rd-order hyperedge $\{A,B,C\}$ with weight 1. On the corresponding quantum state, one can apply different Gaussian operations: (b) Momentum displacement $Z^{B}(-2)$ on mode $B$ adds a hyperedge representing $\operatorname{e}^{-2i\operatorname{\mathit{\hat{q}}}_{B}}$ on it. (c) Position displacement $X^{A}(1)$ on mode $A$ generates new hyperedges: a 1st-order hyperedge representing $\operatorname{e}^{-2i\operatorname{\mathit{\hat{q}}}_{D}}$ on $D$ and a 2nd-order hyperedge representing $\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$ for $\{B,C\}$. (d) Position shearing $D_{q}^{B}(2)$ on mode $B$ adds a decoration representing $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}}$, which is not part of the formalism. (e) Momentum shearing $D_{p}^{A}(1)$ on mode $A$ adds decorations, representing $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$ and $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, which fall outside the formalism, but also induces a valid new hyperedge representing $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$ for $\{B,C,D\}$. (f) Squeezing $S^{A}(1)$ on mode $A$ changes the weights on all hyperedges adjacent to mode $A$. (g) Rotation $R^{D}(\pi)$ on mode $D$ changes the weights by inverting all signs on hyperedges adjacent to $D$. (h) Controlled-$Z$ gate $C_{A,D}(0.5)$ on modes $A$ and $D$ adjusts their weight by 0.5.
@@
In order to present the graphical calculus more conveniently, we first introduce some basic rules of set operations. The hyperedge set $E_{a}$ is defined as the set of all hyperedges containing vertex $a$. The adjacency set $\operatorname{\mathcal{A}}(a)$ of vertex $a\in V$ is defined as $\operatorname{\mathcal{A}}(a)=\{e\setminus\{a\}\mid e\in E_{a}\}$. Considering the bijective mapping between hyperedge and weight sets, we define the union $(E_{U};T_{U})=(E_{1};T_{1})\cup(E_{2};T_{2})$ as a union of all tuples. For hyperedges appearing in both sets, we add up their weights. For the hypergraph in Fig. 2(a), the set of hyperedges containing mode $A$ is given by $E_{A}=\{\{A,B,C\},\{A,D\}\}$ with associated weights $T_{A}=\{1,2\}$ and $\mathcal{A}(A)=\{\{B,C\},\{D\}\}$. If we unite $E_{A}$ and $E_{B}=\{\{A,B,C\}\}$ with $T_{B}=\{1\}$, $(E_{U};T_{U})=(\{\{A,B,C\},\{A,D\}\};\{2,2\})$ is obtained.
@@
Although not all quantum states can be transformed under the graphical rules, some states are able to be represented indirectly in the form of hypergraphs, for example, the state after applying the cubic phase gate $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}$ which contains a third-order moment term of one single-mode operator and plays important roles in universal quantum computing [42, 11, 43]. We will demonstrate later that it can be represented as a sequence of operations similar to those of our formalism.
3 Results
@@
In this section, we provide unified graphical rules of Gaussian unitary operations and Gaussian measurements for CV hypergraph states. We show the rules of single-mode Gaussian operations including displacement, shearing, squeezing, and rotation [44, 45] and further discuss the CV controlled-$Z$ gate. Derivations are given in Appendix A. As any multimode Gaussian operation can be represented as a concatenation of a beam splitter, single-mode squeezing, and phase shifts through the Bloch-Messiah decomposition [46], where the beam-splitter operation is equivalent to a controlled-$Z$ gate [41], any multimode Gaussian operation can be represented in our graphical calculus.
3.1 Displacement
@@
The displacement operators include the momentum displacement $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ and the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$, where $s$ represents the distance that the operator $\operatorname{\mathit{\hat{q}}}_{a}$ or $\operatorname{\mathit{\hat{p}}}_{a}$ is moved in phase space [47]. Thus, the momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ on mode $a$ modifies the hyperedge set by $(E^{\prime};T^{\prime})=(E;T)\cup(\{a\};s)$. In other words, if mode $a$ has its own hyperedge initially, this operation will increase the weight by $s$. If not, a new hyperedge with the weight $s$ will be added to the system. Fig. 2(b) shows an example of performing operation $Z^{B}(-2)$ on the hypergraph state depicted in Fig. 2(a).
@@
However, when performing the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$ to the hypergraph state, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, additional correlation terms will be generated when they simultaneously act on the zero-momentum eigenstate |$\ket{0}_{p}$. It influences hyperedges of the adjacency $\mathcal{A}(a)$ of mode $a$ such that $(E^{\prime},T^{\prime})=(E,T)\cup(\operatorname{\mathcal{A}}(a),\{\tilde{t}\})$, where $\tilde{t}\coloneqq-s\times t_{e\cup a}$, which means the position displacement on mode $a$ will add or change hyperedges between its adjacent modes with the weight multiplied by $-s$. For example, when the position displacement $X^{A}(1)$ is applied to mode $A$ of the hypergraph state corresponding to Fig. 2(a), two new hyperedges between vertices $\{B,C\}$ and $\{D\}$ of $\operatorname{\mathcal{A}}(a)$ are created, as represented by the purple line and circle in Fig. 2(c).
@@
Therefore, in a hypergraph state of $k$-th order, displacement operations can create new hyperedges of at most order $k-1$. Note that for CV graph states, the order of edges is $k=2$ and therefore displacement operations only cause local effects in a certain mode.
3.2 Shearing